A car is traveling to the left, which is the negative direction. The direction of travel remains the same throughout this problem. The car's initial speed is 34.1 m/s, and during a 5.87-second interval, it changes to a final speed of (a)51.1 m/s and (b)30.6 m/s. In each case, find the acceleration (magnitude and algebraic sign)


1
Expert's answer
2022-02-06T14:33:43-0500

Given:

"v_{ix}=-34.1\:\rm m/s"

"t=5.87\:\rm s"


The acceleration of the car

"a_x=\frac{v_{fx}-v_{ix}}{t}"

(a)

"a_x=\frac{-51.1-(-34.1)}{5.87}=-2.90\:\rm m/s^2"

(b)

"a_x=\frac{-30.6-(-34.1)}{5.87}=0.60\:\rm m/s^2"